# A Survey of Secure Semantic Communications

### v1

@@ 4.5.3 Data Generation and Pre-processing
Generative model: Acs et al. [174] introduce differential privacy techniques during the training process of generated models to protect the privacy of the individual data used to train these models. Specifically, the dataset is divided into k clusters using differential privacy and k-mean clustering algorithms. A generative neural network is assigned to each cluster and cluster is trained by differential privacy gradient descent respectively. To solve the problem of complex data privacy protection, Zhang et al. [175] proposed to propose dp-GAN, a generalized private publishing framework for rich semantic data, in which the differential privacy mechanism is integrated based on the existing GAN models. The random noise is injected into the training step of discriminator to perform differential privacy constraints. Various optimization strategies, such as Parameter Grouping, Adaptive Clipping and Warm Starting, are used to generate high-quality synthetic data. Chen et al. [176] introduces a novel approach to secure SemCom by combining differential privacy (DP) with a GAN-based inversion technique to protect sensitive image data transmitted over insecure channels. By extracting and safeguarding essential semantic features, the proposed method reduces privacy risks while enabling accurate data reconstruction at the intended receiver. Simulation results confirm that this approach effectively prevents eavesdroppers from accessing sensitive information while preserving high-quality image reconstruction, highlighting its potential as a robust solution for privacy-preserving SemCom.
@@ 4.6 Cryptography Technology
These emerging encryption technologies offer a balance between data protection and semantic processing. For instance, SMPC enables collaborative data computation without compromising privacy; HE allows addition and multiplication directly on ciphertext, facilitating complex data processing without decryption; TEE ensures secure execution of sensitive data through hardware-level isolation; and Secure Aggregation provides an efficient mechanism for data merging in distributed SemCom. The integration of these technologies not only addresses the shortcomings of traditional encryption but also presents novel solutions for enhancing the efficiency and security of SemCom systems. These new technologies are described below.
@@ 4.6.1 Homomorphic Encryption
PHE: Phong et al. [180] proposed a privacy-preserving DL framework using additive HE. By leveraging the asynchronous stochastic gradient descent algorithm and encrypting gradients with the Paillier system, the framework mitigates potential data leakage risks in cloud-based servers. Experiments showed that this approach achieved near 99% accuracy on the MNIST dataset while effectively preventing data leakage between participants. This research highlights the potential of additive HE in achieving a balance between privacy and performance in distributed learning environments. In FL, Fang et al. [181] introduced a framework combining additive HE and FL, whcih is showed in Fig. 15. This framework incorporates an optimized Paillier algorithm, significantly reducing encryption computation time while maintaining model accuracy deviations within 1%. It is particularly well-suited for applications requiring high privacy standards, such as finance and healthcare. Their experiments further analyzed the impact of key length, network structure, and the number of clients on system performance, providing insights into balancing privacy and efficiency in complex environments.
Figure 15: The architecture of a Paillier federated network [181].
@@
FHE: Lee et al. [182] developed a DL inference model using FHE. Their research employed the CKKS scheme with bootstrapping technology to implement encrypted inference on the ResNet-20 model for the CIFAR-10 dataset, achieving an accuracy of 92.43%, nearly matching the plaintext model’s 91.89%. To address challenges in non-linear activation functions within encrypted environments, they utilized high-precision polynomial approximations as substitutes and, for the first time, implemented Softmax encryption within the FHE framework to defend against model extraction attacks. This study underscores the practicality of FHE in privacy-preserving DL and efficient inference. Additionally, Sun et al. [183] proposed an improved FHE scheme Their work focused on enhancing computational efficiency and reducing noise accumulation through relinearization and modulus switching techniques. This scheme achieved significant efficiency improvements in private decision tree classification and other tasks like Naïve Bayes classification. Experimental results demonstrated its effectiveness, achieving high-performance homomorphic computations across multiple ML classification tasks.
Efficient frameworks and optimized solutions: To address cross-silo FL requirements, Zhang et al. [184] introduced the BatchCrypt system. The system employs a batch encryption method to compress gradient data, significantly reducing encryption computation and communication overhead. Experiments validated that BatchCrypt enhanced training speed by 23–93 times and reduced communication costs by 66–101 times compared to traditional methods, all while maintaining model accuracy. This system provides an efficient and secure encryption framework for cross-organizational data collaboration.
@@ 4.6.2 Secure Multi-Party Computation
Figure 16: The diagram of SMPC [185].
@@ 4.6.3 Trusted Execution Environment
Figure 17: SGX Diagram [195].
Applications in FL: TEE ensures the privacy of training data on local devices. By executing training tasks within a client’s TEE, sensitive data remains in an isolated environment, preventing external access or leakage. For example, Fan et al. [196] introduced a TEE-based privacy protection mechanism in the PPFL framework, which significantly reduces the risks of data reconstruction and inference attacks by locally safeguarding client data and gradients. Additionally, Hamed Haddadi et al. [194] isolate the most sensitive layers of deep neural networks within TEE, effectively enhancing privacy protection. The global model aggregation process in FL is vulnerable to gradient inversion attacks. To address this issue, Yuhui Zhang et al. [195] proposed the ShuffleFL framework, which performs randomization and grouped aggregation within TEE to reduce the likelihood of attackers inferring original data from gradients. On the other hand, Aditya Pribadi Kalapaaking et al. [197] integrated blockchain technology to host model aggregation tasks within TEE, achieving tamper-proof and transparent model updates. Malicious clients in FL may attempt to degrade global model performance by uploading fabricated gradients or reusing outdated model updates. To counter this, Xiaoli Zhang et al. [124] developed the TrustFL framework, utilizing TEE to randomly verify the integrity of client training tasks and ensure adherence to predefined procedures. Furthermore, Arup Mondal et al. [198] proposed FLATEE, which executes model updates in a distributed manner across multiple TEEs, significantly reducing the interference of malicious behavior on aggregation results. TEE can prevent side-channel attacks through hardware isolation and gradient obfuscation techniques. For instance, Yuhui Zhang et al.’s ShuffleFL framework [195] employs a random grouping strategy combined with secure processing within TEE to mitigate attacks through cache and timing analysis. Moreover, the PPFL framework [196] further reduces the risk of gradient leakage by protecting all layers of the model.
@@ 4.6.4 Secure Aggregation
Figure 18: Schematic diagram of the algorithm based on the FFT [200].
@@
Lesson 6
In traditional communication systems, encryption technologies like AES and RSA ensure data confidentiality but limit semantic processing, posing challenges for SemCom. To address these challenges, emerging encryption technologies such as SMPC [187, 188, 189, 190, 191, 192], HE [180, 181, 182, 183, 184], TEE [193, 194, 193, 196, 194, 195], and secure aggregation [199, 200, 201, 202, 203] offer a balance between data protection and semantic processing, enhancing the efficiency and security of SemCom systems. Notably, PHE simplifies operations and is suitable for lightweight privacy-preserving scenarios, while FHE excels in supporting complex tasks with strong security guarantees. With ongoing advancements in algorithms and expanding applications, HE is set to play an increasingly vital role in privacy preservation and collaborative data computation in future SemCom.
@@ 4.7 Block-Chain Technology
Figure 19: Functional diagram of a blockchain network. [212].
@@ 4.7.1 Data Security
Blockchain for sharing semantic information: Lin et al. [215] introduce an innovative framework that integrates blockchain and semantic ecosystems for sharing semantic information. The framework consists of four key components: semantic encoder, channel, semantic proof, and semantic decoder. To address the “garbage-in, garbage-out” problem, the authors propose a semantic consensus mechanism based on threshold signatures, ensuring that only accurate and relevant semantic information is added to the blockchain. Furthermore, a semantic sharing mechanism, utilizing state channels and task-related information bottleneck methods, effectively reduces information redundancy. By combining blockchain’s security and decentralization with the efficient information processing capabilities of SemCom, the framework optimizes network resource utilization while ensuring data security and trust.
@@
Figure 20: An ilustrative framework of Blockchain-Semantic ecosystems [216].
@@ 4.8 Model Compression
Model compression aims at reducing the size of DL models and enhancing the operational efficiency. This is generally achieved by minimizing the model’s computational requirements, such as the number of parameters and computational complexity, without compromising its accuracy significantly. The importance of model compression is paramount when deploying DL models on resource-constrained devices, such as mobile devices, embedded systems, or IoT devices, due to their limited memory, computational power, and battery life. By employing model compression techniques, devices can effectively decrease communication costs, conserve storage space, and eliminate parameter redundancy, thereby strengthening the SemCom system to defend against against multi-party cooperation attacks. Model compression techniques can be divided into the following sub-categories: model pruning, parameter quantization, low-rank decomposition, and knowledge distillation.
@@ 4.8.1 Model Pruning
Figure 21: Combination of model pruning and FL [226].
@@ 4.8.2 Parameter Quantization
Figure 22: Comparison between quantization aware training and post training quantization [231].

### v2

@@ 4.5.3 Data Generation and Pre-processing
Generative model: Acs et al. [169] introduce differential privacy techniques during the training process of generated models to protect the privacy of the individual data used to train these models. Specifically, the dataset is divided into k clusters using differential privacy and k-mean clustering algorithms. A generative neural network is assigned to each cluster and cluster is trained by differential privacy gradient descent respectively. To solve the problem of complex data privacy protection, Zhang et al. [177] proposed to propose dp-GAN, a generalized private publishing framework for rich semantic data, in which the differential privacy mechanism is integrated based on the existing GAN models. The random noise is injected into the training step of discriminator to perform differential privacy constraints. Various optimization strategies, such as Parameter Grouping, Adaptive Clipping and Warm Starting, are used to generate high-quality synthetic data. Chen et al. [178] introduces a novel approach to secure SemCom by combining differential privacy with a GAN-based inversion technique to protect sensitive image data transmitted over insecure channels. By extracting and safeguarding essential semantic features, the proposed method reduces privacy risks while enabling accurate data reconstruction at the intended receiver. Simulation results confirm that this approach effectively prevents eavesdroppers from accessing sensitive information while preserving high-quality image reconstruction, highlighting its potential as a robust solution for privacy-preserving SemCom.
@@ 4.6 Cryptography Technology
These emerging encryption technologies offer a balance between data protection and semantic processing. For instance, SMPC enables collaborative data computation without compromising privacy; HE allows addition and multiplication directly on ciphertext, facilitating complex data processing without decryption; TEE ensures secure execution of sensitive data through hardware-level isolation; and Secure Aggregation provides an efficient mechanism for data merging in distributed SemCom. The integration of these technologies not only addresses the shortcomings of traditional encryption but also presents novel solutions for enhancing the efficiency and security of SemCom systems. These new technologies are described below, and some typical articles are shown in Table LABEL:cryptography.
Table 10: List of Representative Cryptography Schemes
[…]
@@ 4.6.1 Homomorphic Encryption
PHE: Phong et al. [180] proposed a privacy-preserving DL framework using additive HE. By leveraging the asynchronous stochastic gradient descent algorithm and encrypting gradients with the Paillier system, the framework mitigates potential data leakage risks in cloud-based servers. Experiments showed that this approach achieved near 99% accuracy on the MNIST dataset while effectively preventing data leakage between participants. This research highlights the potential of additive HE in achieving a balance between privacy and performance in distributed learning environments. In FL, Fang et al. [181] introduced a framework combining additive HE and FL, whcih is showed in Fig. 15. This framework incorporates an optimized Paillier algorithm, significantly reducing encryption computation time while maintaining model accuracy deviations within 1%. It is particularly well-suited for applications requiring high privacy standards, such as finance and healthcare. Their experiments further analyzed the impact of key length, network structure, and the number of clients on system performance, providing insights into balancing privacy and efficiency in complex environments. To address cross-silo FL requirements, Zhang et al. [192] introduced the BatchCrypt system. The system employs a batch encryption method to compress gradient data, significantly reducing encryption computation and communication overhead. Experiments validated that BatchCrypt enhanced training speed by 23–93 times and reduced communication costs by 66–101 times compared to traditional methods, all while maintaining model accuracy. This system provides an efficient and secure encryption framework for cross-organizational data collaboration.
Figure 15: The architecture of a Paillier federated network [181], where each client requests a key pair from the KMC, initializes the model, and performs forward propagation and backpropagation on encrypted data, ensuring the privacy of sensitive information. During the backpropagation step, the gradients are computed and encrypted before being sent back to the server. The server, which aggregates the gradients from all clients, sends the updated gradients back to each client, facilitating a secure and privacy-preserving training process.
@@
FHE: Lee et al. [182] developed a DL inference model using FHE. Their research employed the CKKS scheme with bootstrapping technology to implement encrypted inference on the ResNet-20 model for the CIFAR-10 dataset, achieving an accuracy of 92.43%, nearly matching the plaintext model’s 91.89%. To address challenges in non-linear activation functions within encrypted environments, they utilized high-precision polynomial approximations as substitutes and, for the first time, implemented Softmax encryption within the FHE framework to defend against model extraction attacks. This study underscores the practicality of FHE in privacy-preserving DL and efficient inference. Additionally, Sun et al. [193] proposed an improved FHE scheme. Their work focused on enhancing computational efficiency and reducing noise accumulation through relinearization and modulus switching techniques. This scheme achieved significant efficiency improvements in private decision tree classification and other tasks like Naïve Bayes classification. Experimental results demonstrated its effectiveness, achieving high-performance homomorphic computations across multiple ML classification tasks.
HE for SemCom: Meng et al. [183] explored the feasibility of FHE in SemCom. They demonstrated that despite traditional encryption techniques disrupting semantic correlation, FHE preserves meaningful features in ciphertext, making it suitable for SemCom applications. Using Scale-Invariant Feature Transform (SIFT), their study validated that encrypted semantic features could still be extracted, ensuring that privacy is maintained while enabling effective communication. Furthermore, they proposed a homomorphic encrypted deep JSCC model, replacing ReLU activation with a square function and using average pooling to support homomorphic operations. Their experiments confirmed that the classification accuracy of their privacy-preserving model on encrypted data was nearly identical to that on plaintext data, demonstrating the potential of FHE in secure SemCom.
@@ 4.6.2 Secure Multi-Party Computation
Figure 16: The diagram of SMPC [194], where multiple parties collaboratively compute a function $f(x_{1},x_{2},\ldots,x_{n})$ with inputs $x_{1},x_{2},\ldots,x_{n}$ and corresponding outputs $y_{1},y_{2},\ldots,y_{n}$ through secure interactions, ensuring data privacy. In this process, each party holds a portion of the data and performs computations on their own inputs, while the final result is shared and computed collaboratively without revealing any individual party’s private input.
@@ 4.6.3 Trusted Execution Environment
Figure 17: SGX Diagram [186], where secret shares and random masks are combined with careful zero padding to reconstruct missing data when some devices drop out, ensuring secure aggregation in federated learning scenarios. The process involves transforming the data using DFT, applying the peeling decoder to extract all shares, and using IDFT to recover the original data, while maintaining data privacy through the use of secret shares and random masks.
Applications in FL: TEE ensures the privacy of training data on local devices. By executing training tasks within a client’s TEE, sensitive data remains in an isolated environment, preventing external access or leakage. For example, Fan et al. [202] introduced a TEE-based privacy protection mechanism in the PPFL framework, which significantly reduces the risks of data reconstruction and inference attacks by locally safeguarding client data and gradients. Additionally, Hamed Haddadi et al. [201] isolate the most sensitive layers of DNNs within TEE, effectively enhancing privacy protection. The global model aggregation process in FL is vulnerable to gradient inversion attacks. To address this issue, Yuhui Zhang et al. [186] proposed the ShuffleFL framework, which performs randomization and grouped aggregation within TEE to reduce the likelihood of attackers inferring original data from gradients. On the other hand, Aditya Pribadi Kalapaaking et al. [203] integrated blockchain technology to host model aggregation tasks within TEE, achieving tamper-proof and transparent model updates. Malicious clients in FL may attempt to degrade global model performance by uploading fabricated gradients or reusing outdated model updates. To counter this, Xiaoli Zhang et al. [125] developed the TrustFL framework, utilizing TEE to randomly verify the integrity of client training tasks and ensure adherence to predefined procedures. Furthermore, Arup Mondal et al. [204] proposed FLATEE, which executes model updates in a distributed manner across multiple TEEs, significantly reducing the interference of malicious behavior on aggregation results. TEE can prevent side-channel attacks through hardware isolation and gradient obfuscation techniques. For instance, Yuhui Zhang et al.’s ShuffleFL framework [186] employs a random grouping strategy combined with secure processing within TEE to mitigate attacks through cache and timing analysis. Moreover, the PPFL framework [202] further reduces the risk of gradient leakage by protecting all layers of the model.
@@ 4.6.4 Secure Aggregation
Figure 18: Schematic diagram of the algorithm based on the FFT [187], where secret shares and random masks are combined with careful zero padding to reconstruct missing data when some devices drop out, ensuring secure aggregation in federated learning scenarios.
@@
4.6.5 Quantum Cryptography
Quantum cryptography leverages the principles of quantum mechanics to provide enhanced security in communication systems. Unlike traditional cryptographic methods that rely on mathematical complexity, quantum cryptography uses the fundamental properties of quantum mechanics, such as superposition and entanglement, to ensure secure communication. This section discusses the integration of quantum cryptography with SemCom, focusing on how quantum techniques can enhance the security and efficiency of semantic data transmission.
Quantum Key Distribution (QKD): One of the most prominent applications of quantum cryptography is QKD. QKD allows two parties to generate a shared random secret key, which can be used for encrypting and decrypting messages. The security of QKD is based on the principles of quantum mechanics, making it immune to computational attacks that can compromise traditional cryptographic keys. In the context of SemCom, QKD can be used to securely transmit semantic keys, ensuring that only authorized parties can access the semantic information. For instance, the study by Chehimi et al. [188] proposes a quantum SemCom framework that integrates QKD to enhance the security of semantic data transmission. The authors demonstrate that QKD can significantly reduce the risk of eavesdropping and ensure the confidentiality of semantic information.
Quantum Digital Signatures(QDS): QDS are another important application of quantum cryptography. QDS provides a way to sign digital messages using quantum states, ensuring the authenticity and integrity of the messages. In SemCom, QDS can be used to sign semantic messages, preventing unauthorized parties from tampering with the semantic information. This is particularly important in scenarios where semantic data needs to be shared among multiple parties, such as in distributed SemCom networks. The work by Nunavath et al. [189] explores the use of QDS in quantum SemCom, highlighting its potential to enhance the security and reliability of semantic data transmission. The authors show that QDS can effectively prevent message tampering and ensure the integrity of semantic information.
Quantum Secure Direct Communication (QSDC): QSDC is a protocol that allows for the direct transmission of secret messages without the need for a pre-shared key. QSDC uses quantum states to encode the messages, ensuring that any eavesdropping attempt will be detected. In SemCom, QSDC can be used to transmit semantic messages directly, providing an additional layer of security. This is particularly useful in scenarios where real-time SemCom is required, such as in emergency response systems or military applications. The study by Khalid et al. [190] investigates the application of QSDC in quantum SemCom, demonstrating its potential to enhance the security and efficiency of semantic data transmission. The authors highlight the advantages of QSDC in terms of reducing latency and improving the reliability of SemCom.
Lesson 5
In traditional communication systems, encryption technologies like AES and RSA ensure data confidentiality but limit semantic processing, posing challenges for SemCom. To address these challenges, emerging encryption technologies such as SMPC [195, 196, 197, 198, 199, 185], HE [180, 181, 182, 193, 192], TEE [200, 201, 200, 202, 201, 186], secure aggregation [205, 187, 206, 207, 208], and Quantum Cryptography [188, 189, 190] offer a balance between data protection and semantic processing, enhancing the efficiency and security of SemCom systems. Notably, PHE simplifies operations and is suitable for lightweight privacy-preserving scenarios, while FHE excels in supporting complex tasks with strong security guarantees. With ongoing advancements in algorithms and expanding applications, HE is set to play an increasingly vital role in privacy preservation and collaborative data computation in future SemCom.
@@ 4.7 Block-Chain Technology
Figure 19: Functional diagram of a blockchain network [217], where blockchain transaction handling starts with an individual initiating a transaction, which is then broadcasted across a decentralized peer-to-peer network of nodes. These nodes are individual computers maintaining a copy of the blockchain ledger, receive the transaction. Miners, specialized nodes equipped with enhanced capabilities, take on the task of verifying the transaction’s legitimacy by checking the sender’s balance, confirming the digital signature, and ensuring the transaction has not been previously spent, thus preventing fraud and maintaining the ledger’s accuracy. Once verified, the miner includes the transaction in a newly created block, which is a collection of transactions, and solves a complex cryptographic puzzle to propose this block to the network.
The below, as shown in Table LABEL:blockchain, the following content, based on multiple academic studies, delves into various blockchain methods for different stages and their potential applications in SemCom.
Table 11: List of Representative Block-Chain Schemes
[…]
@@ 4.7.1 Data Security
Semantic information sharing blockchain framework: Lin et al. [218] propose an innovative framework that combines blockchain technology with semantic ecosystems to facilitate the sharing of semantic information. The framework is composed of four main components: the semantic encoder, channel, semantic proof, and semantic decoder. To tackle the “garbage-in, garbage-out” issue, the authors introduce a semantic consensus mechanism based on threshold signatures, which ensures that only accurate and relevant semantic information is recorded on the blockchain. Additionally, the framework employs a semantic sharing mechanism that utilizes state channels and task-related information bottleneck methods, significantly reducing information redundancy. By leveraging blockchain’s security and decentralization alongside the efficient information processing capabilities of SemCom, the framework enhances network resource utilization while ensuring data security and trust.
@@
Figure 20: An illustrative framework of Blockchain-Semantic ecosystems [218], where begins with raw image data that undergoes semantic encoding to extract meaningful information, which is then transmitted through a wireless channel potentially affected by noise. At the receiving end, a semantic decoder retrieves the information, which is subsequently used for semantic analysis. This analysis involves an aggregated signature that encompasses semantic signatures, semantic strings, and semantic signatures to ensure data integrity and authenticity. Simultaneously, the system operates on a blockchain network where miners validate transactions through semantic verification, ensuring that the information recorded on the blockchain is accurate and tamper-proof. The proof of semantic is crucial here, as it serves as evidence of the semantic integrity of data. Once verified, the semantic information is used to update the blockchain, which is a series of blocks containing verified transactions. The process also includes a state channel that facilitates off-chain transactions for specific tasks, such as image recovery and task-relevant semantic information extraction, enhancing efficiency and reducing load on the main blockchain.
@@ 4.8 Model Compression
Model compression aims at reducing the size of DL models and enhancing the operational efficiency. This is generally achieved by minimizing the model’s computational requirements, such as the number of parameters and computational complexity, without compromising its accuracy significantly. The importance of model compression is paramount when deploying DL models on resource-constrained devices, such as mobile devices, embedded systems, or IoT devices, due to their limited memory, computational power, and battery life. By employing model compression techniques, devices can effectively decrease communication costs, conserve storage space, and eliminate parameter redundancy, thereby strengthening the SemCom system to defend against against multi-party cooperation attacks. Model compression techniques can be divided into the following sub-categories: model pruning, parameter quantization, low-rank decomposition, and knowledge distillation. The representative model compression schemes are listed in Tab. LABEL:modelcompression.
Table 12: List of Representative Model Compression Schemes
[…]
Knowledge Distillation [240] $\Delta$: Proposes the language-oriented SemCom by merging LLMs and generative models with SemCom. $\Theta$: Transmission performance and model complexity. $\Lambda$: Learned Perceptual Image Patch Similarity (LPIPS). $\Xi$: Applied in LLM-aided SemCom for better robustness.
[…]
@@ 4.8.1 Model Pruning
Figure 21: Combination of model pruning and FL [231], where two stages are involved, including initial pruning and further pruning stages. Initial pruning stage: Firstly, it facilitates the commencement of the FL process with a compact model, thereby substantially diminishing the computational and communication overhead per FL round. Secondly, when clients possess varied computational resources, the chosen client for pre-selection pruning can be a powerful and reliable one to expedite the pruning phase. Further pruning stage: The model refined through pre-selection pruning may not be fully optimized, as it is based solely on one client’s data. However, it provides a solid foundation for the FL process encompassing all clients. Throughout the FL process, the standard FedAvg procedure with further adaptive pruning is augmented. This allows the model to dynamically expand or contract, optimizing training efficiency. In this phase, data from all participating clients contribute to the refinement.
@@
Pruning for SemCom: Xie et al. [233] consider an IoT network scenario where the cloud/edge platform handles the training and updating of SemCom models, while IoT devices focus on data collection and transmission based on the trained model. To make it feasible for IoT devices, they propose a lightweight distributed SemCom system for text transmission with low complexity. By pruning the model redundancy and lowering weight resolution, the deep JSCC becomes affordable for IoT devices and significantly decreases the bandwidth required for model weight transmission. Lee et al. [245] introduce a resource-efficient asymmetric autoencoder framework that employs weight pruning and quantization to compress the encoder model on IoT devices. Simulation results show that the encoder-only pruning approach achieves a high compression rate of up to 98% without sacrificing performance.
@@ 4.8.2 Parameter Quantization
Figure 22: Comparison between quantization aware training and post training quantization [246], where (a) Quantization aware training involves taking a pre-trained model, quantizing it, and subsequently fine-tuning it using training data. This fine-tuning helps adjust the parameters and mitigate any accuracy loss. (b): Post training quantization involves calibrating a pre-trained model using calibration data (such as a subset of the training data) to determine clipping ranges and scaling factors. Based on these calibration results, the model is then quantized.
@@
Quantization for SemCom: To facilitate SemComs without end-to-end training, Park et al. [235] introduce a vision transformer-based SemCom system with importance-aware quantization for wireless image transmission. The system leverages the attention scores of a pretrained vision transformer model to assess the importance of image patches. Based on these scores, importance-aware quantization assigns varying quantization bits to patches, prioritizing those with higher importance. This is achieved by solving a weighted quantization error minimization problem, where weights increase with attention scores.